4. A simply supported beam is given below. A movable (crane) concentrated load of 50kips may be placed anywhere along the spaPlease answer and do not copy another solution
Answer #1

Ilive load : 2 kipfft onc load : 50 kips 30 il Draw when influence line diagram unit load is within AB - Efn = 0 I HA =0 IfyDate: Page: alat Mae 30–22) L 30 Mman = 15x15, 30 when amit load is in at 20=153 BD ETKY EMB=0 ta RAX 30 + 1x420 IRR = (1+x)Date: Page:- For A 992 Tallowable = so ksi M : - Goo kelm = 5310.45 kipincha we know M: moment I: moment of inertia si normal